# RegExplainer: Generating Explanations for Graph Neural Networks in Regression Tasks

### v3

@@ Abstract
Graph regression is a fundamental task that has gained significant attention in various graph learning tasks. However, the inference process is often not easily interpretable. Current explanation techniques are limited to understanding GNN behaviors in classification tasks, leaving an explanation gap for graph regression models. In this work, we propose a novel explanation method to interpret the graph regression models (XAIG-R). Our method addresses the distribution shifting problem and continuously ordered decision boundary issues that hinder existing methods away from being applied in regression tasks. We introduce a novel objective based on the graph information bottleneck theory (GIB) and a new mix-up framework, which could support various GNNs and explainers in a model-agnostic manner. Additionally, we present a self-supervised learning strategy to tackle the continuously ordered labels in regression tasks. We evaluate our proposed method on three benchmark datasets and a real-life dataset introduced by us, and extensive experiments demonstrate its effectiveness in interpreting GNN models in regression tasks.
@@ 1 Introduction
However, existing methods focus on the explanation of the classification tasks, leaving another fundamental task, explainable regression, unexplored. Graph regression tasks exist widely in nowadays applications, such as predicting the molecular property [17] or traffic flow volume [18]. Explaining the instance-level predictions of graph regression is challenging due to two main obstacles. First, in the routinely adopted GIB framework, the mutual information between the explanation sub-graph and label, $I(G^{*};Y)$, is estimated with the Cross-Entropy between the predictions $f(G^{*})$ from GNN model $f$ and its prediction label $Y$. However, in the regression task, the regression label is the continuous value, making the approximation unsuitable. Another challenge is the distribution shifting problem in the usage of $f(G^{*})$, the prediction of the explanation sub-graph made by the GNN model $f$. Usually, explanation sub-graphs have different topology and feature information compared to the original graph. As a result, explanation sub-graphs are out-of-distribution of the original training graph dataset. As shown in Figure 1, a GNN model $f$ is trained on the original graph training set and cannot be safely used to make predictions for sub-graphs.
@@
$\bullet$ We proposed a novel model with self-supervised learning and the mix-up approach, which could more effectively address the two challenges, and better explain the graph model on the regression tasks compared to other baselines.
$\bullet$ We designed three synthetic datasets, namely BA-Motif-Volume, BA-Motif-Counting and Triangles, as well as a real-world dataset called Crippen, which could also be used in future works, to evaluate the effectiveness of our regression task explanations. Comprehensive empirical studies on both synthetic and real-world datasets demonstrate that our method can provide consistent and concise explanations for graph regression tasks.
@@ 2 Related Work and Further Discussions
The explanation methods for GNN models could be categorized into two types based on their granularity: instance-level [19, 20, 21, 22] and model-level [23], where the former methods explain the prediction for each instance by identifying important sub-graphs, and the latter method aims to understand the global decision rules captured by the GNN. These methods could also be classified into two categories based on their methodology: self-explainable GNNs [24, 25] and post-hoc explanation methods [20, 21, 22], where the former methods provide both predictions and explanations, while the latter methods use an additional model or strategy to explain the target GNN. Additionally, CGE [26] (cooperative explanation) generates the sub-graph explanation with the sub-network simultaneously, by using cooperative learning. However, it has to treat the GNN model as a white box, which is usually unavailable in the post-hoc explanation. Existing methods haven’t explored the explanation of the graph regression task and haven’t considered two important challenges: the distribution shifting problem and the limitation of the GIB objective, which were addressed by our work.
@@ 4.1.1 Optimizing the Lower Bound of $I(G^{*};Y)$
Intuitively, the property of $I(Y^{*};Y)$ is guaranteed by the chain rule for mutual information and the independence between each explanation instance $g^{*}$. An intuitive demonstration is shown in Figure 2. The proof is shown in the Appendix B.1.
@@ 4.1.2 Estimating $I(Y^{*};Y)$ with InfoNCE
Now the challenge becomes the estimation of the mutual information $I(Y^{*};Y)$. Inspired by the model of Contrastive Predictive Coding [33], in which InfoNCE loss is interpreted as a mutual information estimator, we further adapt the objective function so that it could be applied with InfoNCE loss in explaining graph regression. In our graph explanation scenario, the InfoNCE Loss defined in Eq. (3) could also be utilized as a lower bound of $I(Y^{*};Y)$, as shown in the following property with proofs:
@@
where $Y^{\prime}$ is the prediction label of the graph randomly sampled graph neighbors, $\mathbb{Y}$ is the set of the neighbors’ prediction labels, and $\text{sim}()$ estimates the similarity between $Y^{*}$ and $Y$. The proof is shown in the Appendix B.2. Therefore, we have the InfoNCE loss $\mathcal{L}_{\text{NCE}}$ as the lower bound of the $I(Y^{*};Y)$. We approximate Eq. (2) as:
@@ 4.2 Distribution Shifting Problem in Graph Regression
We include the prediction label $Y^{*}$ in Eq. (4) to estimate similarity, which is approximated with $f(G^{*})$ in previous work [20, 21]. However, we argue that $f(G^{*})$ cannot be safely obtained due to the distribution shift problem [34, 35]. In classification tasks, a small shift may not cross the decision boundaries, which could still lead to a correct prediction. However, due to the continuous decision boundaries in regression, the distribution problem would cause serious prediction errors. Here in this paper, the graph distribution is indicated by its regression label in the regression task.
@@
Intuitively, this phenomenon indicates the GNN model $f$ could make correct predictions only with the original graph $G$ yet could not predict the explanation sub-graph $G^{*}$ correctly. This is because the GNN model $f$ is trained with the original graph sets, whereas the explanation $G^{*}$ as the sub-graph is different from the original graph sets. With the shift between $f(G)$ and $f(G^{*})$, the optimal solution in Eq. (4) is unlikely to work well.
@@ 4.3 Mix-up Approach with Contrastive Learning
$\bullet$ Step 1 (Neighbor Sampling): Learning through the triplet instances could effectively reinforce the ability of the explainer to learn the explanation self-supervised. For each target graph $G$ with label $Y$ to be explained, we can define two randomly sampled graphs as positive neighbor $G^{+}$ and negative neighbor $G^{-}$, where $G^{+}$’s label $Y^{+}$ is closer to $Y$ than $G^{-}$’s label $Y^{-}$, i.e., $|Y^{+}-Y|<|Y^{-}-Y|$. Intuitively, the distance between the distributions of the positive pair $\langle G,G^{+}\rangle$ should be smaller than the distance between the distributions of the negative pair $\langle G,G^{-}\rangle$.
@@
$G^{\text{(mix)}+}$ and $G^{\text{(mix)}-}$ should have the similar information to $G$ because they have the same label-preserving sub-graphs $G^{*}$. Additionally, considering the following two pairs: $(G^{\text{(mix)}+},G^{+})$ and $(G^{\text{(mix)}-},G^{-})$. The similarity between $(G^{\text{(mix)}+},G^{+})$ should be larger than the similarity between $(G^{\text{(mix)}-},G^{-})$. Intuitively, since $G^{\text{(mix)}+}$ and $G^{\text{(mix)}-}$ have the same label-preserving sub-graphs $G^{*}$ and $|Y^{-}-Y|>|Y^{+}-Y|$, we could have $|f(G^{-})-f(G^{\text{(mix)}-})|>|f(G^{+})-f(G^{\text{(mix)}+})|$, where $f(G)$ represents the prediction label of graph $G$.
$\bullet$ Step 3 (InfoNCE Loss Approximation): Then we could safely estimate the similarity with $\text{sim}(Y^{\text{(mix)}},Y)$. To save more information, we use the similarity of representation embedding to approximate the similarity of the graph prediction label, where ${\bm{h}}^{\text{(mix)}}$ represents the embedding for $G^{\text{(mix)}}$ and ${\bm{h}}$ represents the embedding for $G$. We use $\mathbb{H}$ to represent the neighbors set accordingly. Thus, we approximate Eq. (4) as:
@@
The mix-up approach in previous work [35] generates a mixed graph by simply mixing explanation sub-graph $G^{*}$ with a randomly sampled label-irrelevant sub-graph $G^{\Delta}$ [35], which could be formally written as $G_{a}^{\text{(mix)}}=G_{a}^{*}+(G_{b}-G_{b}^{*})$. However, it couldn’t tackle the continuous decision boundaries in graph regression tasks. A detailed description of the mix-up approach can be found in Appendix C.
@@ 4.4 Implementation
After generating the mix-up explanation $G^{\text{(mix)}}$, we specify the InfoNCE loss to further train the parameterized explainer with a triplet of graphs $\langle G,G^{+},G^{-}\rangle$. In practice, $G^{+}$ and $G^{-}$ are randomly sampled from the graph dataset, upon which we calculate their similarity score with the target graph $G$. The sample with a higher score would be the positive sample and the other one would be the negative sample. Specifically, we use $\text{sim}({\bm{h}},{\bm{h}}^{\prime})={\bm{h}}^{\top}{\bm{h}}^{\prime}$ to compute the similarity score, where $G^{\prime}$ could be $G^{+}$ or $G^{-}$. ${\bm{h}}$ is generated by feeding $G$ into the GNN model $f$ and retrieving the embedding vector before the dense layers.
@@
In practice, the denominator in Eq. (5) works as a regularization to avoid trivial solutions. Since the label $Y$ is given and independent of the optimization process, we could also employ the MSE loss between $Y^{*}$ and $Y$ additionally, regarding InfoNCE loss only estimates the mutual information between the embeddings. Formally, the overall loss function could be implemented as:
@@ 5.1 Experiment Settings
We formulate Three synthetic datasets and a real-world dataset, as is shown in Table 1, in order to address the lack of graph regression datasets with ground-truth explanation. The datasets include: BA-Motif-Volume and BA-Motif-Counting, which are based on BA-shapes [20], Triangles [37], and Crippen [38]. We compared the proposed RegExplainer against a comprehensive set of baselines in all datasets, including: GRAD  [20], ATT  [39], GNNExplainer [20], PGExplainer [21], and MixupExplainer [35]. Detailed information about experiment setting are put in the Appendix E. We elaborate on the measurement metric of methods as follows: (1) AUC-ROC: We use the AUC score to evaluate the performance of our proposed methods against baseline methods regarding the ground-truth explanation, which could be treated as a binary classification task. (2) We evaluate the similarity of distribution of the graph with Cosine Similarity and Euclidean Distance.
@@ 5.4 Alleviating Distribution Shifts (RQ3)
We observed that in Figure 3, red points surround the blue points but green points are shifted away, which indicates that the explanation sub-graph couldn’t help GNNs make correct predictions. As shown in Table 2, we calculate the RMSE score between the $f(G)$ and $Y$, $f(G^{*})$ and $Y$, $f(G)$ and $f(G^{*})$ respectively, where $f(G)$ is the prediction the original graph, $f(G^{*}$ is the prediction of the explanation sub-graph, and $Y$ is the regression label. We could observe that $f(G^{*})$ shows a significant prediction shifting from $f(G)$ and $Y$, indicating that the mutual information calculated with $(f(G^{*}),Y)$ would be biased.
@@
We also design experiments to illustrate how RegExplainer corrects the deviations: we calculate the graph embeddings $v$ and predictions $p$ of the explanation sub-graphs and the mix-up graph. Then we compare them to the ground truth and calculate the Euclidean or Cosine distance between the vectors and RMSE between prediction labels. From the results in Table 3, we can observe that all the performances of COS($v_{g}$, $v_{m}$), EUC($v_{g}$, $v_{m}$) and prediction errors are better than those of ($v_{g}$, $v_{e}$), which indicates RegExplainer could effectively fix the distribution of sub-graph explanation $G^{*}$ and reduce the embedding distance and prediction error.
@@ 6 Conclusion
We addressed the challenges in the explainability of graph regression tasks and proposed the RegExplainer, a novel method for explaining the predictions of GNNs with the post-hoc explanation sub-graph on graph regression task without requiring modification of the underlying GNN architecture or re-training. We showed how RegExplainer can leverage the mix-up approach to solve the distribution shifting problem and adopt the GIB objective with the InfoNCE loss to migrate it from graph classification tasks to graph regression tasks, while these existing challenges seriously affect the performances of other explainers. We formulated four new datasets: BA-Motif-Volume, BA-Motif-Counting, Triangles, and Crippen for evaluating the explainers on the graph regression task, which are aligned with the design of datasets in previous work. They could also benefit future studies on the XAIG-R. While we acknowledge the effectiveness of our method, we also recognize its limitations. Specifically, although our approach can be applied to explainers for graph regression tasks in an explainer-agnostic manner, it cannot be easily applied to explainers built for explaining the spatio-temporal graph due to the dynamic topology and node features of the STG. To overcome this challenge, a potential solution is to incorporate cached dynamic embedding memories into the framework.

### v4

@@ Abstract
Graph regression is a fundamental task that has gained significant attention in various graph learning tasks. However, the inference process is often not easily interpretable. Current explanation techniques are limited to understanding Graph Neural Network (GNN) behaviors in classification tasks, leaving an explanation gap for graph regression models. In this work, we propose a novel explanation method to interpret the graph regression models (XAIG-R). Our method addresses the distribution shifting problem and continuously ordered decision boundary issues that hinder existing methods away from being applied in regression tasks. We introduce a novel objective based on the graph information bottleneck theory (GIB) and a new mix-up framework, which can support various GNNs and explainers in a model-agnostic manner. Additionally, we present a self-supervised learning strategy to tackle the continuously ordered labels in regression tasks. We evaluate our proposed method on three benchmark datasets and a real-life dataset introduced by us, and extensive experiments demonstrate its effectiveness in interpreting GNN models in regression tasks.
@@ 1 Introduction
However, existing methods focus on the explanation of the classification tasks, leaving another fundamental task, explainable regression, unexplored. Graph regression tasks exist widely in nowadays applications, such as predicting the molecular property [17] or traffic flow volume [18]. Explaining the instance-level predictions of graph regression is challenging due to two main obstacles. First, in the routinely adopted GIB framework, the mutual information between the explanation sub-graph and label, $I(G^{*};Y)$, is estimated with the Cross-Entropy between the predictions $f(G^{*})$ from GNN model $f$ and its prediction label $Y$. However, in the regression task, the regression label is the continuous value, making the approximation unsuitable. Another challenge is the distribution shifting problem in the usage of $f(G^{*})$, where the prediction of the explanation sub-graph made by the GNN model $f$ is unsafe. Usually, explanation sub-graphs have different topology and feature information compared to the original graph. As a result, explanation sub-graphs are out-of-distribution of the original training graph dataset [19, 20, 21]. As shown in Figure 1, a GNN model $f$ is trained on the original graph training set and cannot be safely used to make predictions for sub-graphs.
@@
$\bullet$ We proposed a novel model with self-supervised learning and the mix-up approach, which can address the two challenges more effectively, and better explain the graph model on the regression tasks compared to other baselines.
$\bullet$ We designed three synthetic datasets, namely BA-Motif-Volume, BA-Motif-Counting and Triangles, as well as a real-world dataset called Crippen, which can also be used in future works, to evaluate the effectiveness of our regression task explanations. Comprehensive empirical studies on both synthetic and real-world datasets demonstrate that our method can provide consistent and concise explanations for graph regression tasks.
@@ 2 Related Work and Further Discussions
The explanation methods for GNN models can be categorized into two types based on their granularity: instance-level [22, 23, 24, 25] and model-level [26], where the former methods explain the prediction for each instance by identifying important sub-graphs, and the latter method aims to understand the global decision rules captured by the GNN. These methods can also be classified into two categories based on their methodology: self-explainable GNNs [27, 28] and post-hoc explanation methods [23, 24, 25], where the former methods provide both predictions and explanations, while the latter methods use an additional model or strategy to explain the target GNN. Additionally, CGE [29] (cooperative explanation) generates the sub-graph explanation with the sub-network simultaneously, by using cooperative learning. However, it has to treat the GNN model as a white box, which is usually unavailable in the post-hoc explanation. Existing methods have only partially addressed the explanation of graph regression tasks and have not fully considered two important challenges: the distribution shifting problem and the limitations of the GIB objective, both of which are key areas our work aims to tackle.
@@ 4.1.1 Optimizing the Lower Bound of $I(G^{*};Y)$
Intuitively, the property of $I(Y^{*};Y)$ is guaranteed by the chain rule for mutual information and the independence between each explanation instance $g^{*}$ in $G^{*}$. An intuitive demonstration is shown in Figure 2. The proof is shown in the Appendix B.1.
@@ 4.1.2 Estimating $I(Y^{*};Y)$ with InfoNCE
Now the challenge becomes the estimation of the mutual information $I(Y^{*};Y)$. Inspired by the model of Contrastive Predictive Coding [36], in which InfoNCE loss is interpreted as a mutual information estimator, we further adapt the objective function so that it can be applied with InfoNCE loss in explaining graph regression. In our graph explanation scenario, the InfoNCE Loss defined in Eq. (3) can also be utilized as a lower bound of $I(Y^{*};Y)$, as shown in the following property with proofs:
@@
where $Y^{\prime}$ is the prediction label of the randomly sampled graph neighbors, $\mathbb{Y}$ is the set of the neighbors’ prediction labels, and $\text{sim}()$ estimates the similarity between $Y^{*}$ and $Y$. The proof is shown in the Appendix B.2. Therefore, we have the InfoNCE loss $\mathcal{L}_{\text{NCE}}$ as the lower bound of the $I(Y^{*};Y)$. We approximate Eq. (2) as:
@@ 4.2 Distribution Shifting Problem in Graph Regression
We include the prediction label $Y^{*}$ in Eq. (4) to estimate similarity, which is approximated with $f(G^{*})$ in previous work [23, 24]. However, we argue that $f(G^{*})$ cannot be safely obtained due to the distribution shift problem [37, 19]. In classification tasks, a small shift may not cross the decision boundaries, which can still lead to a correct prediction. However, due to the continuous decision boundaries in regression, the distribution problem would cause serious prediction errors. Here in this paper, the graph distribution is indicated by its regression label in the regression task.
@@
Intuitively, this phenomenon indicates the GNN model $f$ can make correct predictions only with the original graph $G$ yet can not predict the explanation sub-graph $G^{*}$ correctly. This is because the GNN model $f$ is trained with the original graph sets, whereas the explanation $G^{*}$ as the sub-graph is different from the original graph sets. With the shift between $f(G)$ and $f(G^{*})$, the optimal solution in Eq. (4) is unlikely to work well.
@@ 4.3 Mix-up Approach with Contrastive Learning
$\bullet$ Step 1 (Neighbor Sampling): Learning through the triplet instances can effectively reinforce the ability of the explainer to learn the explanation self-supervised. For each target graph $G$ with label $Y$ to be explained, we can define two randomly sampled graphs as positive neighbor $G^{+}$ and negative neighbor $G^{-}$, where $G^{+}$’s label $Y^{+}$ is closer to $Y$ than $G^{-}$’s label $Y^{-}$, i.e., $|Y^{+}-Y|<|Y^{-}-Y|$. Intuitively, the distance between the distributions of the positive pair $\langle G,G^{+}\rangle$ should be smaller than the distance between the distributions of the negative pair $\langle G,G^{-}\rangle$.
@@
$G^{\text{(mix)}+}$ and $G^{\text{(mix)}-}$ should have the similar information to $G$ because they have the same label-preserving sub-graphs $G^{*}$. Additionally, considering the following two pairs: $(G^{\text{(mix)}+},G^{+})$ and $(G^{\text{(mix)}-},G^{-})$. The similarity between $(G^{\text{(mix)}+},G^{+})$ should be larger than the similarity between $(G^{\text{(mix)}-},G^{-})$. Intuitively, since $G^{\text{(mix)}+}$ and $G^{\text{(mix)}-}$ have the same label-preserving sub-graphs $G^{*}$ and $|Y^{-}-Y|>|Y^{+}-Y|$, we can have $|f(G^{-})-f(G^{\text{(mix)}-})|>|f(G^{+})-f(G^{\text{(mix)}+})|$, where $f(G)$ represents the prediction label of graph $G$.
$\bullet$ Step 3 (InfoNCE Loss Approximation): Then we can safely estimate the similarity with $\text{sim}(Y^{\text{(mix)}},Y)$. To save more information, we use the similarity of representation embedding to approximate the similarity of the graph prediction label, where ${\bm{h}}^{\text{(mix)}}$ represents the embedding for $G^{\text{(mix)}}$ and ${\bm{h}}$ represents the embedding for $G$. We use $\mathbb{H}$ to represent the neighbors set accordingly. Thus, we approximate Eq. (4) as:
@@
The mix-up approach in previous work [19] generates a mixed graph by simply mixing explanation sub-graph $G^{*}$ with a randomly sampled label-irrelevant sub-graph $G^{\Delta}$ [19], which can be formally written as $G_{a}^{\text{(mix)}}=G_{a}^{*}+(G_{b}-G_{b}^{*})$. However, it cann’t tackle the continuous decision boundaries in graph regression tasks. A detailed description of the mix-up approach can be found in Appendix C.
@@ 4.4 Implementation
After generating the mix-up explanation $G^{\text{(mix)}}$, we specify the InfoNCE loss to further train the parameterized explainer with a triplet of graphs $\langle G,G^{+},G^{-}\rangle$. In practice, $G^{+}$ and $G^{-}$ are randomly sampled from the graph dataset, upon which we calculate their similarity score with the target graph $G$. The sample with a higher score would be the positive sample and the other one would be the negative sample. Specifically, we use $\text{sim}({\bm{h}},{\bm{h}}^{\prime})={\bm{h}}^{\top}{\bm{h}}^{\prime}$ to compute the similarity score, where $G^{\prime}$ can be $G^{+}$ or $G^{-}$. ${\bm{h}}$ is generated by feeding $G$ into the GNN model $f$ and retrieving the embedding vector before the dense layers.
@@
In practice, the denominator in Eq. (5) works as a regularization to avoid trivial solutions. Since the label $Y$ is given and independent of the optimization process, we can also employ the MSE loss between $Y^{*}$ and $Y$ additionally, regarding InfoNCE loss only estimates the mutual information between the embeddings. Formally, the overall loss function can be implemented as:
@@ 5.1 Experiment Settings
We formulate Three synthetic datasets and a real-world dataset, as is shown in Table 1, in order to address the lack of graph regression datasets with ground-truth explanation. The datasets include: BA-Motif-Volume and BA-Motif-Counting, which are based on BA-shapes [23], Triangles [39], and Crippen [40]. We compared the proposed RegExplainer against a comprehensive set of baselines in all datasets, including: GRAD  [23], ATT  [41], GNNExplainer [23], PGExplainer [24], and MixupExplainer [19]. Detailed information about experiment setting are put in the Appendix E. We elaborate on the measurement metric of methods as follows: (1) AUC-ROC: We use the AUC score to evaluate the performance of our proposed methods against baseline methods regarding the ground-truth explanation, which can be treated as a binary classification task. (2) We evaluate the similarity of distribution of the graph with Cosine Similarity and Euclidean Distance.
@@ 5.4 Alleviating Distribution Shifts (RQ3)
We observed that in Figure 3, red points surround the blue points but green points are shifted away, which indicates that the explanation sub-graph cann’t help GNNs make correct predictions. As shown in Table 2, we calculate the RMSE score between the $f(G)$ and $Y$, $f(G^{*})$ and $Y$, $f(G)$ and $f(G^{*})$ respectively, where $f(G)$ is the prediction the original graph, $f(G^{*}$ is the prediction of the explanation sub-graph, and $Y$ is the regression label. We can observe that $f(G^{*})$ shows a significant prediction shifting from $f(G)$ and $Y$, indicating that the mutual information calculated with $(f(G^{*}),Y)$ would be biased.
@@
We also design experiments to illustrate how RegExplainer corrects the deviations: we calculate the graph embeddings $v$ and predictions $p$ of the explanation sub-graphs and the mix-up graph. Then we compare them to the ground truth and calculate the Euclidean or Cosine distance between the vectors and RMSE between prediction labels. From the results in Table 3, we can observe that all the performances of COS($v_{g}$, $v_{m}$), EUC($v_{g}$, $v_{m}$) and prediction errors are better than those of ($v_{g}$, $v_{e}$), which indicates RegExplainer can effectively fix the distribution of sub-graph explanation $G^{*}$ and reduce the embedding distance and prediction error.
@@ 6 Conclusion
We addressed the challenges in the explainability of graph regression tasks and proposed the RegExplainer, a novel method for explaining the predictions of GNNs with the post-hoc explanation sub-graph on graph regression task without requiring modification of the underlying GNN architecture or re-training. We showed how RegExplainer can leverage the mix-up approach to solve the distribution shifting problem and adopt the GIB objective with the InfoNCE loss to migrate it from graph classification tasks to graph regression tasks, while these existing challenges seriously affect the performances of other explainers. We formulated four new datasets: BA-Motif-Volume, BA-Motif-Counting, Triangles, and Crippen for evaluating the explainers on the graph regression task, which are aligned with the design of datasets in previous work. They can also benefit future studies on the XAIG-R. While we acknowledge the effectiveness of our method, we also recognize its limitations. Specifically, although our approach can be applied to explainers for graph regression tasks in an explainer-agnostic manner, it cannot be easily applied to explainers built for explaining the spatio-temporal graph due to the dynamic topology and node features of the STG. To overcome this challenge, a potential solution is to incorporate cached dynamic embedding memories into the framework.